OpenStudy (loser66):

Solve the system\[x'=\left[\begin{matrix}1&1&0\\1&1&0\\0&0&3\end{matrix}\right]x+\left[\begin{matrix}e^t\\e^{2t}\\3e^{3t}\end{matrix}\right]\] Anyone walk me through, please

OpenStudy (loser66):

any way of U.C, Diagonalize, Variation parameter or Laplace, please

OpenStudy (loser66):

@oldrin.bataku @SithsAndGiggles @ybarrap @amistre64

OpenStudy (anonymous):

I would use UC, but only because I've forgotten how to apply the other methods. Homogeneous solution: \[x'=\left[\begin{matrix}1&1&0\\1&1&0\\0&0&3\end{matrix}\right]x\] \[\begin{vmatrix}1-\lambda&1&0\\1&1-\lambda&0\\0&0&3-\lambda\end{vmatrix}=0~\Rightarrow~\lambda_1=0,\lambda_2=2,\lambda_3=3\] Which gives you the following corresponding eigenvectors: \[\eta_1=\begin{pmatrix}1\\-1\\0\end{pmatrix},\eta_2=\begin{pmatrix}1\\1\\0\end{pmatrix},\eta_3=\begin{pmatrix}0\\0\\1\end{pmatrix}\] So you have the homogeneous solution of \[y_c=C_1\eta_1e^{\lambda_1 t}+C_2\eta_2e^{\lambda_2 t}+C_3\eta_3e^{\lambda_3 t}\\ y_c=C_1\eta_1+C_2\eta_2e^{2 t}+C_3\eta_3e^{3 t}\]

OpenStudy (anonymous):

Sorry, meant \(x_c\) instead of \(y_c\), but you get the idea. As a guess for the nonhomogeneous solution, I'd try \[x_p=e^t\vec{a}+te^{2t}\vec{b}+te^{3t}\vec{c}\] I'm not entirely sure if that would work, but I haven't tried it yet.

OpenStudy (anonymous):

I use that guess because you can write the last vector as three vectors: \[e^t\begin{pmatrix}1\\0\\0\end{pmatrix}+e^{2t}\begin{pmatrix}0\\1\\0\end{pmatrix}+e^{3t}\begin{pmatrix}0\\0\\1\end{pmatrix}\]